OpenStudy (anonymous):

What wavelength of light is needed to eject an electron from an aluminum atom given its binding energy is 9.60 x 10-19 J?

OpenStudy (aaronq):

E= hc/lambda

OpenStudy (anonymous):

can you show me how to do that

OpenStudy (aaronq):

plug it in, and solve like algebra E=hc/lambda lambda = hc/E lambda= hc/9.60 x 10-19 J h = plancks constant, c = speed of light

OpenStudy (anonymous):

6.626*10^-34*3.00*10^8 is that it

OpenStudy (anonymous):

/9.60*10^-19

OpenStudy (aaronq):

yep